OpenStudy (anonymous):

When n is even, integrals of the form ∫tan^m(x)sec^n(x) dx can be evaluated by factoring out sec^2(x)=1+tan^2(x) and using the fact that Dx tanx=sec^2(x). When m is odd, integrals of this form can be evaluated by factoring out tanxsecx and using the fact that Dx secx=secxtanx. Use this method to evaluate the following integral: ∫tan^(-3/2)(x) sec^4(x) dx

zepdrix (zepdrix):

Is this for calc 2 or something? Man you have some tricky problems! :O lol \[\large \int\limits \frac{\sec^4x}{\tan^{3/2}x}dx\] So `n is even`, (the power on the secant). Ok ok ok that's good.

OpenStudy (anonymous):

yes calc 2 unfortunately

zepdrix (zepdrix):

We'll turn 2 of the powers on secant into tangents.\[\large \int\limits\limits \frac{\sec^2x\sec^2x}{\tan^{3/2}x}dx \qquad = \qquad \large \int\limits\limits \frac{(1+\tan^2x)\sec^2x}{\tan^{3/2}x}dx\]

zepdrix (zepdrix):

It might be easier to see what's going on if we move stuff around a little.\[\large \int\limits \frac{1+\tan^2x}{\tan^{3/2}x}(\sec^2x\;dx)\]See how we have a bunch of ugly tangent stuff... and thennnnn we also have the derivative of tangent sitting in those brackets!!! :O Is the U sub jumping out at you yet?? :D

OpenStudy (anonymous):

ok, let me try to see what i get

zepdrix (zepdrix):

Oh man I hated calc 2... These types of problems were actually a lot of fun in my opinions. I love all kinds of integrals. Area between curves, trig subs.. all that fun stuff :) Partial fractions, long division, impropers, fun fun fun. Power series is where it got very difficult for me :C

OpenStudy (anonymous):

lol ok im having trouble with this

OpenStudy (anonymous):

and yes, not liking calc 2, but need it for my major. hopefully i can learn to love integrals like you though :)

zepdrix (zepdrix):

\[\large u=\tan x\]\[\large du=\sec^2x\;dx\] \[\large \int\limits\limits \frac{1+\tan^2x}{\tan^{3/2}x}(\sec^2x\;dx) \qquad \rightarrow \qquad \large \int\limits\limits \frac{1+u^2}{u^{3/2}}(du)\] Were you able to get this far?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

From here we'll split up the fraction,\[\large \int\limits\frac{1+u^2}{u^{3/2}}du \qquad = \qquad \int\limits \frac{1}{u^{3/2}}+\frac{u^2}{u^{3/2}}du\]Divide the terms, apply negative exponents so it's easier to read, and integrate! :)

OpenStudy (anonymous):

ok will give it a try now

OpenStudy (anonymous):

so then we get 2(u^2-3)/3u^(1/2)

OpenStudy (anonymous):

is that right so far?

zepdrix (zepdrix):

Hmmm I'm not sureeeeee what's going on there, why don't you have two terms? Are you being all fancy and combining them? XD lol

zepdrix (zepdrix):

\[\large \int\limits\limits u^{-3/2}+u^{1/2}\;du\]Then apply the power rule to each term.

OpenStudy (anonymous):

i got it...so I put tanx into my u and i got [2(tan^2(x)-3)]/[3sqrt tanx]

OpenStudy (anonymous):

lol sorry...just started playing around with it...but i definitely should separate them in future ...

OpenStudy (anonymous):

thanks so much for your help zepdrix :)

zepdrix (zepdrix):

yay team \c:/